A goods train completely crosses a 120 m long bridge with uniform speed 'v' in 't' seconds. Then it will
cross 240 m long bridge with uniform speed 2v in.

Grade:12th Pass

2 Answers

vikas askiitian expert
13 years ago

T = distance/speed =  120/v    (initial time)

final time = 240/2v = 120/v     (final time)

therefore time will remain same ......

Ritvik Gautam
13 years ago

We know,

s=ut+1/2 at2

but since the train is going with uniform velocity, hence the acceleration is zero.

therefore,

s=ut

In the first case,

120=vt

t=120/v

In the second case,

240=2vt

t=120/v

 

Hence, the time in both the cases, time taken will be the same.

 

Please Approve!!